If cosec (theta) = root 10 & theta is an acute angle, then sec (theta) = ?
Found 2 solutions by Alan3354, Fombitz:
Answer by Alan3354(69443)

If cosec (theta) = root 10 & theta is an acute angle, then sec (theta) = ?
--------------
sin = sqrt(10)/10
sin^2 + cos^2 = 1
---
cos^2 = 1 - sin^2 = 1 - 1/10 = 9/10
sec^2 = 1/cos^2 = 10/9
sec = sqrt(10)/3

Answer by Fombitz(32388)